OpenStudy (lizzymonster):

The area of a rectangle is 65 ft2, and the length of the rectangle is 3 ft less than twice the width. Find the dimensions of the rectangle.

OpenStudy (lizzymonster):

What I have so far: 65ft^2=w(2w-3) 65ft^2=2w^2-3w 2w^2-3w-65 and then you're supposed to factor into two binomials, but I cant seem to find any factors that multiply to equal 65 and add make -3

Vocaloid (vocaloid):

2w^2-3w-65 factors to (w+5)(2w-13)

OpenStudy (lizzymonster):

but -13+5 doesnt equal -3, so how do you know?

Vocaloid (vocaloid):

the rule works a bit differently because we have (2w^2) instead of just w^2 we can prove my factoring is correct using FOIL (w+5)(2w-13) = (2w)(w) - 13(w) + 10w - 65 = 2w^2 - 3w - 65

Vocaloid (vocaloid):

this video may be helpful https://www.youtube.com/watch?v=07g-l8OuGPo

OpenStudy (lizzymonster):

ok thank you!